# HMD2: Environment-aware Motion Generation from Single Egocentric Head-Mounted Device

### v1

@@ 4.1 Main Results
Quantitative Results.  The main quantitative results are summarized in Table 1, with a finer-grained analysis based on different activity categories provided in the supplementary. Our system achieved superior performance across all three metric axes of reconstruction, diversity, and realism. The online variant of our system degrades performance slightly, as expected, given its inaccessibility of future sensor information.
Our adapted version of AvatarPoser (referred to as AvatarPoser (Head) in Table 1) performs well, but its frame-by-frame prediction can lack temporal coherence, reducing realism. As a regression model, it captures only the average trend in training data, leading to lower diversity scores. Unlike our multi-modal model, AvatarPoser doesn’t incorporate environmental context, which enhances our system’s performance (Fig. 4). EgoEgo also generates visually reasonable motions but has two key issues. Despite its Diffusion-based long motion inference, discontinuities in long motions affect realism metrics. Additionally, EgoEgo tends to produce overly dynamic arm movements, similar to how some image Diffusion models create stylized rather than naturalistic outputs. Using all repetitions instead of the lowest-error repetition for metrics, this leads to higher Hand Position Errors and contributes to increased MPJPE and Diversity scores compared to ground truth. Note that while all the metrics are measured as mean across all runs, we additionally report MPJPE of the best-case run: 8.246, 14.678, for HMD2 and EgoEgo (AvatarPoser is deterministic, so stays the same). Compared to Table 1, errors for EgoEgo are noticeably lower but are still behind Ours.
@@
Table 2: Ablation study. HMD2 leverages both point cloud (PC) and ego-centric video information (CLIP) to reduce per-joint error while keeping the realism and physical plausibility of the motions.
@@ 5 Conclusions and Discussions
We presented a Diffusion-based framework, HMD2, for online motion generation from a single head-mounted device. By leveraging camera streams for learning-based image embeddings and combining them with SLAM-derived head trajectories and semi-dense point clouds, our framework can produce diverse and natural motions that align with environmental contexts. We assessed our system across various environments and an extensive range of daily activities. Compared to existing state-of-the-art methods, HMD2 significantly enhances motion quality in terms of accuracy, diversity, and realism.
@@ A Technical details
Our conditional motion diffusion model follows the Transformer-based architecture presented in EDGE [63] with additional MLP encoder layers to gradually reduce the input dimension (which is bigger due to added CLIP and PC features) to the token latent space size. Our input consists of the motion input (as a translation, rotation, and linear and angular velocities) and PC and CLIP features, all concatenated together, representing one sequence token per frame. Following AvatarPoser [32], the model only predicts local joint rotations but not global translation. The global movement of the character is created during test time by “stitching” the predicted body motion to the ground-truth head motion, and the head motion can be directly obtained through real HMD motion obtained through SLAM, offset by a constant calibration matrix provided by the dataset. The motion output of the EDGE Diffusion model is denoted as $\bm{x}\in\mathbb{R}^{T\times F}$, where $T=240$ and $F=23\times 6$. The skeleton following Xsens definition has 23 ball-and-socket joints, and for each joint, the output rotation is represented as the first two columns of its local rotation matrix. Note that the definition of Xsens human skeleton is very similar to SMPL [42], with the main difference being the ordering of joints. The model is not conditioned on body size information, but during training, it is forced to see HMD input motions from different subjects covering highly diverse demographics. As such, the trained model is able to handle body size variation implicitly. However, providing size information as an explicit condition might further improve model performance and reduce visual artifacts such as floor penetration and foot sliding. To create the motion visualizations and compute position error metrics, we used ground truth body sizes (skeleton bone lengths) for each subject.
@@
System runtime.  Our current implementation assumes that point cloud encodings and CLIP features are precomputed or computed in parallel on a separate device. The performance will be affected if all computations need to happen on the same device. However, we observed that even in this situation, we could achieve a throughput of $\sim 61$ FPS for our low-latency variant, therefore keeping up with realtime speed: CLIP embeddings take around 5 ms to compute per image (2.5 ms per motion frame since we are duplicating every frame), and point cloud encoder taking around 0.1 ms per motion frame. Note that the runtime performance is evaluated on a powerful GPU, which indicates a gap for our system to work in realtime on board of the HMD itself. Additionally, our current implementation assumes the access of all SLAM feature points in the around 15min window of the whole motion sequence. In a true realtime setting, this simplification would require a warm-up phase in the same environment of similar time length.
@@ B Dataset details
We split the dataset for training/validation/testing as 806/10/224 sequences, corresponding to 202/3/56 hours. The testing split does not contain any locations or subjects that appear in the training set to ensure no data leakage. We also strive to maintain a similar distribution of activities between the training set and the test set. This dataset will be released in recent months.

### v2

@@ 4.1 Main Results
Quantitative Results.  The main quantitative results are summarized in Table 1, with a finer-grained analysis provided in the supplementary. Our system achieved superior performance across all three metric axes of reconstruction, diversity, and realism. As expected, the online variant of our system degrades performance slightly, given inaccessibility of future sensor information, but still outperforms baselines.
Our adapted version of AvatarPoser (referred to as AvatarPoser (Head) in Table 1) performs well, but its frame-by-frame prediction lacks temporal coherence, reducing realism. As a regression model, it captures only the average trend in training data, leading to lower diversity scores. Unlike our multi-modal approach, it lacks environmental awareness, impacting performance (Fig. 4). EgoEgo generates plausible motions but has two key issues. First, it produces discontinuities during long motion inference, which affect realism metrics. Second, EgoEgo tends to produce overly dynamic arm movements, similar to how some image diffusion models create stylized rather than naturalistic outputs. This leads to higher Hand Position Errors and contributes to increased MPJPE and Diversity scores compared to ground truth. While all the metrics in Table 1 are measured as mean across all runs, we additionally report MPJPE of the best-case run: 8.246, 14.678, for HMD2 and EgoEgo (AvatarPoser stays the same). Compared to Table 1, errors for EgoEgo are noticeably lower but are still behind Ours.
@@
Table 2: Ablation study. HMD2 leverages both point cloud (PC) and egocentric video information (CLIP) to reduce per-joint error while keeping the realism and physical plausibility of the motions.
@@ 5 Conclusions and Discussions
We introduced HMD2, a diffusion-based framework for online motion generation from a single head-mounted device. By combining camera-based image embeddings with SLAM-derived head trajectories and semi-dense point clouds, HMD2 produces diverse, natural motions aligned with the environment. Our evaluation across various settings and activities shows that HMD2 outperforms state-of-the-art methods in accuracy, diversity, and realism.
@@ A Technical details
Our conditional motion diffusion model follows the Transformer-based architectures presented in EDGE [63] and DiT [53] with additional MLP encoder layers to gradually reduce the input dimension (which is bigger due to added CLIP and PC features) to the token latent space size. Our input consists of the motion input (as a translation, rotation, and linear and angular velocities) and PC and CLIP features, all concatenated together, representing one sequence token per frame. Following AvatarPoser [32], the model only predicts local joint rotations but not global translation. The global movement of the character is created during test time by “stitching” the predicted body motion to the ground-truth head motion, and the head motion can be directly obtained through real HMD motion obtained through SLAM, offset by a constant calibration matrix provided by the dataset. The motion output of the diffusion model is denoted as $\bm{x}\in\mathbb{R}^{T\times F}$, where $T=240$ and $F=23\times 6$. The skeleton following Xsens definition has 23 ball-and-socket joints, and for each joint, the output rotation is represented as the first two columns of its local rotation matrix. Note that the definition of Xsens human skeleton is very similar to SMPL [42], with the main difference being the ordering of joints. The model is not conditioned on body size information, but during training, it is forced to see HMD input motions from different subjects covering highly diverse demographics. As such, the trained model is able to handle body size variation implicitly. However, providing size information as an explicit condition might further improve model performance and reduce visual artifacts such as floor penetration and foot sliding. To create the motion visualizations and compute position error metrics, we used ground truth body sizes (skeleton bone lengths) for each subject.
@@
System runtime.  Our current implementation assumes that point cloud encodings and CLIP features are precomputed or computed in parallel on a separate device. The performance will be affected if all computations need to happen on the same device. However, we observed that even in this situation, we could achieve a throughput of $\sim 61$ FPS for our low-latency variant, therefore keeping up with real-time speed: CLIP embeddings take around 5 ms to compute per image (2.5 ms per motion frame since we are duplicating every frame), and point cloud encoder taking around 0.1 ms per motion frame. Note that the runtime performance is evaluated on a powerful GPU, which indicates a gap for our system to work in real-time on board of the HMD itself. Additionally, our current implementation assumes the access of all SLAM feature points in the around 15min window of the whole motion sequence. In a true real-time setting, this simplification would require a warm-up phase in the same environment of similar time length.
@@ B Dataset details
We split the dataset for training/validation/testing as 806/10/224 sequences, corresponding to 202/3/56 hours. The testing split does not contain any locations or subjects that appear in the training set to ensure no data leakage. We also strive to maintain a similar distribution of activities between the training set and the test set.